If f(x, y) = 1, then the value of ∫∫R 1 dxdy gives,,, choose the correct answer?


A) 2 * Area of the region R


B) 1/2* Area of the region

C) Area of the region R R


D) 0


1
Expert's answer
2021-09-03T01:59:45-0400

A) 2 * Area of the region R


A derivative is a function that measures the slope. It depends upon x in some way and is found by differentiating a form y = f (x) function.  When x is substituted into the derivative, the result is the original function y = f (x) slope.

There are many different ways to indicate the operation of differentiation, also known as finding or taking the derivative. The choice of notation depends on the type of function being evaluated and upon personal preference.

Suppose you have a general function: y = f(x). All of the following notations can be read as "the derivative of y concerning x" or, less formally, "the derivative of the function."